the odds of macy passing her driver's test on the first try are 12:13. Determine the probability that she will pass her driver's test
a) 0.923 b) 0.520 c) 0.480 d) 1.083

Answer by jim_thompson5910(35256)


"the odds of macy passing her driver's test on the first try are 12:13" means that Macy will pass 12 times out of 12+13 = 25 total.
The ratio 12:13 indicates there are 12 successes for every 13 failures. This is assuming "odds in favor".
For "odds against", you would flip the ratio. Eg: Saying "odds against a horse winning a race is 20:1" means there are 20 failures for every 1 success

If we assume "odds in favor", then divide the two values (12 and 25) to get: 12/25 = 48/100 = 0.480